### v2

# WebCiteS: Attributed Query-Focused Summarization on Chinese Web Search Results with Citations

Abstract
Enhancing the attribution in large language models (LLMs) is a crucial task. One feasible approach is to enable LLMs to cite external sources that support their generations. However, existing datasets and evaluation methods in this domain still exhibit notable limitations. In this work, we formulate the task of attributed query-focused summarization (AQFS) and present WebCiteS, a Chinese dataset featuring 7k human-annotated summaries with citations. WebCiteS derives from real-world user queries and web search results, offering a valuable resource for model training and evaluation. Prior works in attribution evaluation do not differentiate between groundedness errors and citation errors. They also fall short in automatically verifying sentences that draw partial support from multiple sources. We tackle these issues by developing detailed metrics and enabling the automatic evaluator to decompose the sentences into sub-claims for fine-grained verification. Our comprehensive evaluation of both open-source and proprietary models on WebCiteS highlights the challenge LLMs face in correctly citing sources, underscoring the necessity for further improvement. [^1] The dataset and code are released under Apache License 2.0 at https://github.com/HarlynDN/WebCiteS
[…]
WebCiteS: Attributed Query-Focused Summarization on Chinese Web Search Results with Citations
Haolin Deng1111thanks: This work was done during the internship at Tencent. Chang Wang2222  Xin Li2222  Dezhang Yuan2222  Junlang Zhan2222 Tianhua Zhou2222Jin Ma3333Jun Gao111122footnotemark: 2Ruifeng Xu1111thanks: Corresponding authors $1$Harbin Institute of Technology, Shenzhen, China $2$Tencent $3$University of Science and Technology of China hldeng028@gmail.com, dezhangyuan@tencent.com, xuruifeng@hit.edu.cn
1 Introduction
In today’s information-driven society, swift access to knowledge is essential. A major limitation of web search engines is the need for users to manually compile information from various sources, which can be time-consuming. Large language models (LLMs) Zhao et al. (2023) exhibit potential in this domain by generating straightforward and well-organized responses. However, the potential risks of hallucination (Ji et al., 2023; Zhang et al., 2023c) and factual errors (Min et al., 2023) undermine their trustworthiness as knowledge sources.
An emerging solution is generative search engines (Liu et al., 2023a) which use LLMs to synthesize web search results into responses with in-line citations. This allows users and developers to verify the generations against the cited sources. However, recent investigations on commercial products and retrieval-augmented LLMs reveal frequent occurrences of unsupported claims and incorrect citations (Liu et al., 2023a; Gao et al., 2023b), highlighting the challenges of attribution in LLMs (Li et al., 2023a). Nonetheless, the limitations of pertinent datasets and evaluation methods pose obstacles to in-depth explorations within the community.
Figure 1: Illustration of attributed query-focused summarization (AQFS). Full example is shown in Table 10.
Lan.
Query
Source
Document
Source
# Cnt.
Docs
Length
Response
Length
# Citations
per Sent.
Citation
Annot.
Citation
Eval.
WebCiteS ZH Real user queries
Web search
results
7k 1025 / 3970 167 1.55 human auto
WebCPM
(Qin et al., 2023)
ZH
Translated
Reddit questions
Web searh
results
5k 513 244 0.34 human N/A
WebGLM
Liu et al. (2023b)
EN ELI5
Web searh
results
45k 304 104 1.20 auto human
ALCE
(Gao et al., 2023b)
EN
ASQA
QAMPARI, ELI5
Wikipedia
Common Crawl
3k 100*k 78 N/A N/A auto
[…]
Firstly, most existing datasets are deficient in high-quality citation annotations.
For instance, the ALCE benchmark (Gao et al., 2023b) compiles three question-answering datasets (Fan et al., 2019; Stelmakh et al., 2022; Amouyal et al., 2023) without providing citations in the reference answers, limiting its utility for model training. In contrast, WebGLM  (Liu et al., 2023b) prompts InstructGPT (Ouyang et al., 2022) to generate training data with citations. It controls the citation quality via a sample filtering method which calculates the ROUGE Lin (2004) score between the answers and their citations. However, this method focuses on lexical similarity rather than logical entailment, and thus could not precisely measure attribution.
Secondly, current evaluation methods are insufficient to thoroughly assess attribution.
Prior works only inspect if the generations are supported by their citations Liu et al. (2023a); Gao et al. (2023b); Liu et al. (2023b) without checking all the documents provided in the input context. However, instances of unsupported generations may result from both failing to correctly cite supporting documents and failing to be grounded in all input documents. Differentiating these two types of errors is crucial for system optimization. Moreover, existing automatic evaluation (Gao et al., 2023b) solely relies on off-the-shelf natural language inference (NLI) models which only recognize entailment (full support) and overlook sentences with multiple sub-claims drawing partial support from different sources. Such complexities are common in real-world scenarios (Chen et al., 2023; Kamoi et al., 2023) and are indicative of a strong capability of synthesizing information across various sources.
To address the above limitations, we present WebCiteS, a Chinese dataset for Attributed Query-Focused Summarization (AQFS). As shown in Figure 1, given a query and the retrieved documents, AQFS aims to summarize all pertinent information from the documents with in-line citations to make the summary attributable. Our dataset is built upon real-world user queries and search results from Sogou, a widely used Chinese web search engine.222www.sogou.com We employ human efforts to ensure the quality of summaries and citations. Table 1 compares WebCiteS and relevant datasets.
We propose a comprehensive evaluation framework with a cost-effective automatic evaluator. Our evaluation metrics distinguish two key aspects: groundedness (if the model outputs are contextually supported) and citation quality (citation accuracy and comprehensiveness), enabling a more nuanced understanding of attribution errors. We also train a tailored claim-split model to extract the sub-claims of a sentence for fine-grained verification. This allows the detection of partial support and improves the alignment between our automatic evaluator and human citations.
Our evaluation of both open-source and proprietary models on WebCiteS reveals the following key findings: (1) contextual grounding of generations does not guarantee the avoidance of citation errors, indicating the challenge of explicit attribution in all the tested LLMs; (2) although supervised fine-tuning improves both groundedness and citation quality, the top-performing model only reaches a citation F1 score of 76.1% with about 20% of sentences not being fully supported by their citations, underscoring the need for further optimization; (3) models perform worse when summarizing full content of web pages rather than shorter snippets, showing that LLMs are less effective at synthesizing and attributing information in the longer context; (4) making documents more fine-grained leads to poorer attribution results, highlighting the difficulty LLMs face in pinpointing the exact supporting evidence within the context.
Figure 2: Illustration of the human-LLM collaborative annotation pipeline of WebCiteS. Initially, annotators manually extract useful information from the documents; then, LLMs are used to generate candidate summaries from the extraction; finally, annotators choose the preferred candidate, refine its quality, and annotate citations.
2 The WebCiteS Dataset
In this section, we first formulate the AQFS task and then delineate the construction of WebCiteS.
2.1 Task Formulation of AQFS
For a query $q$ and its set of retrieved documents $\mathcal{D}$, AQFS aims to generate a summary $\mathcal{S}$. Following previous works Liu et al. (2023a); Gao et al. (2023b), we segment it into sentences: $\mathcal{S}=\{s_{1},\ldots,s_{n}\}$. Each sentence $s_{i}$ cites a subset of documents $\mathcal{C}_{i}=\{d_{1},d_{2},\ldots\}$ where $d_{i}\in\mathcal{D}$. Citations are only required for sentences deemed verification-worthy Liu et al. (2023a), i.e., sentences needing external evidence for validation. We formulate this property with citation mask =$\mathcal{M}=\{m_{1},...,m_{n}\}$, where $m_{i}$ is a binary value and $m_{i}=1$ denotes sentence $s_{i}$ requires citations. In practice, we find most of the sentences require citations, except the introductory or concluding sentences in the summary, such as “There are several possible reasons for papaya tasting a bit bitter”.
2.2 Data Collection
We collected 40,000 unique, anonymized user queries from Sogou, a widely used Chinese web search engine, encapsulating a diverse range of real-world questions. After initial refinement, we retained 18,500 non-trivial queries and retrieved five web pages for each query.333 Common trivial queries include ones that look for word synonyms, text translation, celebrity birth dates, etc. During annotation, we found the top five search results were adequate to address most queries. The snippets of web pages returned by the search engine were expanded to a maximum of 250 characters to serve as the documents for the annotation process.
2.3 Human-LLM Collaborative Annotation
Crafting a comprehensive long-form summary from various documents is challenging and labor-intensive for human annotators. Meanwhile, LLMs have showcased impressive proficiency in certain annotation tasks He et al. (2023). With this in mind, we conceptualized a collaborative annotation pipeline of three stages, as illustrated in Figure 2.
Stage 1: Manual Screening and Information Extraction.
Firstly, human annotators read the query and documents thoroughly. They would extract all useful information from the documents and evaluate its utility. We found over 95% of the queries could be answered by the extracted information, and a few exceptions were discarded.
Stage 2: LLM-based Candidate Summary Generation.
We leveraged LLMs to construct candidate summaries from the extracted information. Generating summaries from human-extracted content, as opposed to raw documents, avoided the introduction of irrelevant information. We employed ChatGPT [^4] We use the gpt-3.5-turbo-0613 checkpoint in this stage. in the preliminary annotation phase. As our dataset grew, we fine-tuned an open-source model, ChatGLM2-6B Du et al. (2022), to provide an extra candidate summary for each sample.
Stage 3: Manual Refinement and Citation Annotation.
Lastly, human annotators chose the preferred summary among the two LLM-generated candidates, refined its quality, and annotated citations. The chosen summary underwent thorough inspection with non-essential and redundant parts removed. Annotators would cite all supporting documents for verification-worthy sentences, correct groundedness and coherence errors, and supplement missing information. Offering multiple candidate summaries aimed to avoid limiting annotators to a single, potentially lower-quality option and to merge the strengths of different options, thereby improving the quality of the final summary.
Statistic Value
# Train / Dev / test 5,630 / 500 / 1,000
# Domains 16
# Search Results per Query 5
Full Content Len. / Snippet Len. 774 / 205
Summary Len. 167
# Sent. per Summary 4.56
# Citations / Sub-claims per Sent. 1.55 / 1.62
Table 2: Core statistics of the WebCiteS dataset. Full content length and snippet length refer to the average length of a single search result (web page). Sub-claims of sentences are extracted by ChatGPT (Section 4.2).
Quality control.
We collaborated with crowd-sourcing companies for data labeling. We recruited a team of 27 annotators, 7 quality inspectors, and 1 senior quality inspector, all underwent a month-long training. The quality inspectors reviewed all annotations from the first and third stages, and the senior inspector randomly checked the passed ones. Annotations that failed to meet the standards were returned for corrections and re-inspected.
The core statistics of WebCiteS are present in Table 2. Moreover, we conduct the following analysis on the quality of the dataset:
Are the retrieved documents useful to the queries?
Though we did not ask annotators to explicitly label the relevance score for each document, the human-extracted information from stage 1 could reflect how useful the document is to the query. After annotation, we find that 87.2% of the documents have human extraction, while the average length of extracted segments per document is 93.4 characters. This indicates that the majority of retrieved documents are helpful to the query.
How much manual refinement is made on candidate summaries?
The average Levenshtein distance between the human-refined summary and the human-preferred candidate summary is 74.1 (we only count summary edits, excluding citation annotation). This suggests that the quality of candidate summaries were generally judged imperfect by the human annotators.
Overlap of web pages.
We also notice that previous works Krishna et al. (2021); Lewis et al. (2021); Bolotova et al. (2023) point out a high test-train overlap within the dataset would hinder the models to ground generations in the context. While the queries in WebCiteS are non-repetitive, we additionally examine the URLs of all searched web pages serving as documents in different data splits, and find only 0.3% of the URLs in the train splits exist in validation and test splits, eliminating the concern of high test-train overlap.
See Appendix A for more details of the dataset.
3 Evaluation Framework
The AQFS task targets two dimensions: summarization utility and attribution. In this section, we introduce the evaluation metrics based on an evaluator with two key components: a claim-split model $\psi$ and an NLI model $\phi$. The claim-split model decomposes each sentence $s_{i}$ into a set of sub-claims $\psi(s_{i})=\{c_{i,1},c_{i,2},\ldots\}$. The NLI model $\phi$ predicts if the given premise entails, contradicts, or is neutral to the given hypothesis. We report the performance of $\phi$ and $\psi$ in Section 4.
3.1 Evaluating Summarization Utility
We adopt the following metrics to evaluate the utility of the summaries:
Length.
We report the average summary length, as prior studies (Gehrmann et al., 2023; Liu et al., 2023c) point out that the summary lengths across different systems exhibit a large variance which is not well-reflected by other metrics.
Self-BLEU.
Self-BLEU (Zhu et al., 2018) measures the diversity of the generated text. Xu et al. (2023) find this metric effective at evaluating the coherence of long-form answers.
Claim precision.
We apply $\psi$ to extract all sub-claims from the system summary and calculate the fraction of them being entailed by the reference summary using $\phi$. This metric could measure the accuracy and relevance of system summaries.
Claim recall.
Similarly, we apply $\psi$ to extract all sub-claims from the reference summary and calculate the fraction of them being entailed by the system summary. This metric could measure the comprehensiveness of system summaries.
Claim $\text{F}_{1}$.
Finally, we can compute the claim $\text{F}_{1}$ score of a system by taking the harmonic mean of its claim precision and recall. See Appendix B.1 for more discussions on summarization metrics.
3.2 Evaluating Attribution
Only verification-worthy sentences with citation mask $m_{i}=1$ are included in attribution evaluation. Gao et al.’s (2023b) automatic metrics assume that all sentences need citations, i.e. the citation mask $m_{i}$ is always 1. However, we observe some exceptions such as the introductory or concluding sentences in the summaries (see Section 2.1). Therefore, we propose to automatically predict the citation mask for sentence $s_{i}$:
$m_{i}=\mathds{1}(\mathcal{C}_{i}\neq\emptyset\vee\phi(\mathcal{S}^{*},s_{i})\neq\text{entailment})$
[…]
For sentences with citation mask $m_{i}=1$, we evaluate two aspects for attribution: groundedness and citation quality with the following metrics:
AIS.
The AIS score assesses if the generation is attributable to identified sources (Bohnet et al., 2022; Rashkin et al., 2023). We adopt the fine-grained version proposed in RARR (Gao et al., 2023a), which calculates the fraction of attributable sentences in the generation. Since citations serve as the identified sources in AQFS, a sentence $s_{i}$ is attributable if it is fully supported by its citations $\mathcal{C}_{i}$. In practice, our automatic evaluator will classify $s_{i}$ as attributable if (1) $s_{i}$ does not contradict any citation in $\mathcal{C}_{i}$ and (2) $s_{i}$ or all of its sub-claims $\psi(s_{i})$ are entailed by the citations. Under this definition, the AIS score is equivalent to the citation recall metric proposed in ALCE (Gao et al., 2023b). Since this metric generally measures both citation quality and groundedness, we adopt the term AIS and define a variant of citation recall.
ACS.
We propose attributable to contextual sources (ACS), a variant of the AIS score that uses oracle citations from the evaluator rather than the model-generated citations for evaluation. This isolates groundedness assessment by eliminating the impact of citation errors. For example, if $s_{i}$ is contextually grounded but does not cite any source, its AIS and ACS scores will be 0 and 1 respectively.
Citation precision.
This metric measures if the sentence is correctly supported by each of its citations. For $s_{i}$, we extract the model-generated citations $C^{i}_{\text{pred}}$ and obtain the oracle citations $C^{i}_{\text{ref}}$ involving all $d_{j}\in\mathcal{D}$ that fully or partially support $s_{i}$. Then, citation precision for $s_{i}$ is:
Citation $\text{Citation Prec}(s_{i})=\frac{\lvert C_{\text{pred}}^{i}\cap C_{\text{ref}}^{i}\rvert}{\lvert C_{\text{pred}}^{i}\rvert}$
In practice, the evaluator would include $d_{j}$ into $C^{i}_{\text{ref}}$ if (1) it entails $s_{i}$ or (2) it does not contradict $s_{i}$ and entails any of its sub-claims. Moreover, if $C^{i}_{\text{pred}}$ is empty, we seek the nearest non-empty citations $C^{i*}_{\text{Pred}}$ from the subsequent sentences of $s_{i}$ to replace $C_{\text{Pred}}^{i}$ for evaluation. This is to avoid penalizing the model for generating multiple sentences based on the same sources but only adding citations in the final sentence. If $C^{i*}_{\text{Pred}}$ is not found, citation precision for $s_{i}$ woule be zero. Finally, we average the scores of all $s_{i}\in\mathcal{S},m_{i}=1$ as the citation precision score for the summary $\mathcal{S}$.
Citation recall.
This metric measures if the sentence comprehensively cites all supporting sources:
Citation $\text{Citation Rec}(s_{i})=\frac{\lvert C_{\text{pred}}^{i}\cap C_{\text{ref}}^{i}\rvert}{\lvert C_{\text{ref}}^{i}\rvert}$
Similarly, we seek $C^{i*}_{\text{Pred}}$ if $C^{i}_{\text{pred}}$ is empty. We assign a zero score if $C^{i}_{\text{Ref}}$ is empty. Finally, we average the citation recall scores of all $s_{i}\in\mathcal{S},m_{i}=1$.
Citation $\text{F}_{1}$.
Similar to Claim $\text{F}_{1}$, we compute the citation $\text{F}_{1}$ score of a system by taking the harmonic mean of its citation precision and recall.
Figure 3 shows the framework of attribution evaluation. Compared to Gao et al.’s (2023b) automatic methods, our method (1) considers citation mask, (2) incorporates the claim-split model for partial support detection, and (3) distinguishes groundedness and citation quality. See Appendix B.2 for more discussions on attribution metrics.
Figure 3: Illustration of our attribution evaluation. We use a claim-split model to extract sub-claims of a sentence and conduct fine-grained verification on all the source documents. The translation is in italic text.
4 Evaluating the Automatic Evaluator
In this section, we assess the reliability of the automatic evaluator with $\phi$ and $\psi$.
4.1 Performance of the NLI model
We evaluate the performance of different open-source NLI models in predicting human-annotated citations in WebCiteS. We finally choose an mT5 model (Xue et al., 2021) fine-tuned on multilingual NLI tasks as $\phi$ for our evaluator,666huggingface.co/alan-turing-institute/mt5-large-finetuned-mnli-xtreme-xnli since it achieves the highest accuracy of 82.3% among all models. See Appendix C for more details about NLI models.
4.2 Performance of the Claim-Split Model
We first prompt ChatGPT [^7] We use the gpt-3.5-turbo-1106 checkpoint. to extract the sub-claims in sentences since Kamoi et al. (2023) have validated this approach via human assessment. As using proprietary LLMs for automatic evaluation still faces limitations in efficiency and cost, we additionally fine-tune mT5 on the outputs of ChatGPT to learn this task. We evaluate the claim-split models with the following metrics:
Redundancy.
It measures if the model generates redundant sub-claims of the source sentence. Two sub-claims $c_{i}$ and $c_{j}$ in $\psi(s)$ are deemed redundant if they entail each other: $\phi(c_{i},c_{j})=1$ and $\phi(c_{j},c_{i})=1$. Based on this, we could eliminate the redundancy of $\psi(s)$: if multiple sub-claims are redundant, we only keep the first one and remove the rest. The resulting set is denoted as $\psi^{*}(s)$. Finally, the metric is computed as:
$\text{Redundancy}(\psi(s))=\frac{\rvert\psi(s)\rvert-\rvert\psi^{*}(s)\rvert}{\rvert\psi(s)\rvert}$
# Splits.
It is defined as the average count of non-redundant sub-claims per sentence, which could reflect the granularity of model outputs, as a lower value indicates that the model may not effectively separate some sub-claims in the source sentence.
Correctness.
It is defined as the fraction of sub-claims being entailed by the source sentence using the NLI model $\phi$, as we assume that all correct sub-claims should be entailed by the source sentence.
Completeness.
It is a binary value measuring if the source sentence is entailed by the concatenation of all sub-claims using the NLI model $\phi$. If not, some essential sub-claims may be missing in the model outputs.
Table 3 shows the evaluation result. We first notice that ChatGPT and the fine-tuned models are consistent in # splits which reflects the decomposition granularity. Moreover, the fine-tuned models slightly outperform ChatGPT in correctness and completeness. The primary reason is that ChatGPT occasionally unfollow the task instruction (Zhou et al., 2023). Therefore, we select the fine-tuned mT5-large model as $\psi$ for our evaluator. See Appendix D for more details about claim-split models.
Redun. # Splits Correct. Complete.
mT5-Base 2.4 1.9 99.0 90.0
mT5-Large 1.7 1.9 99.0 92.2
ChatGPT 1.7 1.9 96.7 89.3
Table 3: Performance of different claim-split models.
Claim-Split
Citation
Prec.
Citation
Rec.
AIS
AIS
Default Citation Mask
-- 77.7 74.5 84.8 90.3
mT5-Large 77.7 82.3 87.3 93.1
ChatGPT 77.9 81.0 87.1 92.7
Auto Citation Mask
-- 81.7 73.4 84.8 89.8
mT5-Large 81.7 82.3 87.4 92.7
ChatGPT 81.9 80.7 87.1 92.4
Human Citation Mask
-- 82.4 74.0 85.4 90.0
mT5-Large 82.4 83.0 88.2 93.1
ChatGPT 82.6 81.3 87.9 92.7
Table 4: Performance of the automatic evaluator on evaluating the attribution in human-annotated summaries. We use a fixed NLI model and vary the use of claim-split models under different citation mask settings.
4.3 Performance of the Automatic Evaluator
Finally, we assess our automatic evaluator with $\phi$ and $\psi$ on the test set of WebCiteS. We use it to predict the citations for sentences in the reference summaries, and then assess those citations by taking human citations as ground truth. We also compute the AIS scores using both citations. We use the same NLI model $\phi$ and vary the use of $\psi$ to analyze the impact of the claim-split strategy. Besides, we compare different citation mask settings: (1) default, which sets $m_{i}=1$ for all sentences; (2) auto, which automatically predicts $m_{i}$ (see Section 3.2); (3) human, which only sets $m_{i}=1$ for sentences with human citations. The results in Table 4 show that: (1) Claim-split model helps to detect partial support. Integrating $\psi$ enhances both citation recall and AIS scores, indicating that the citations that only support part of the sub-claims of the sentences are effectively identified; (2) Accurate citation mask improves the performance of the evaluator. Using the human citation mask yields the best overall performance, while the auto citation mask achieves better citation precision compared to the default setting. This result emphasizes the necessity of identifying if a sentence requires citations during evaluation.
Moreover, the Cohen’s kappa coefficient between the evaluator (using mT5-large as $\psi$ and auto citation mask) and human annotators on whether a sentence should cite a document is 0.6483, which suggests substantial agreement. This further validates the reliability of the automatic evaluator.
Len. Self- Claim Citation AIS ACS
BLEU Prec. Rec. F1 Prec. Rec. F1
FSP ChatGPT 223 12.2 53.5 56.5 54.9 71.2 64.8 67.8 75.1 84.7
GPT-4 285 5.6 46.1 65.8 54.2 72.2 71.4 71.8 75.7 81.1
ChatGLM2-6B 256 11.7 40.1 45.1 42.5 9.9 7.3 8.4 9.8 73.4
ChatGLM3-6B 252 10.0 45.6 53.7 49.3 57.0 51.4 54.1 60.0 85.0
Baichuan2-7B 188 7.6 46.9 48.4 47.6 16.3 21.6 18.6 20.8 81.0
Baichuan2-13B 200 7.8 45.4 49.6 47.4 44.4 51.0 47.4 52.3 76.2
SFT mT5-Base 142 19.7 47.2 35.9 40.8 47.2 35.9 37.7 40.0 82.0
mT5-Large 142 15.7 52.9 40.6 45.9 50.2 51.1 50.6 56.0 90.3
ChatGLM2-6B 156 8.7 54.6 47.6 50.9 76.3 72.5 74.4 79.4 87.1
ChatGLM3-6B 163 8.9 55.7 49.3 52.3 78.5 73.8 76.1 81.3 89.4
Baichuan2-7B 212 9.9 52.4 53.1 52.8 68.1 67.8 67.9 71.6 81.7
Baichuan2-13B 132 5.3 58.9 42.1 49.1 67.7 72.0 69.8 74.4 81.0
Table 5: Results of the AQFS task on WebCiteS, where each sample consists of five snippets as documents.
5 Experiments on the AQFS Task
We evaluate various models on the AQFS task via two methods: few-shot prompting (FSP) and supervised fine-tuning (SFT). Our prompt for FSP consists of the task instruction and one-shot demonstration, while the SFT prompt removes the demonstration and condenses the instruction for efficiency. For open-source models, we evaluate mT5, ChatGLM2-6B, ChatGLM3-6B (Du et al., 2022), Baichuan2-7B, and Baichuan2-13B (Yang et al., 2023) via both FSP and SFT. [^8] We use the chat version of Baichuan2 models. For proprietary models, we evaluate ChatGPT and GPT-4 via FSP. [^9] We use the gpt-3.5-turbo-1106 checkpoint for ChatGPT and the gpt-4-1106-preview checkpoint for GPT-4. See Appendix E.1 for implementation details.
5.1 Main Results
We first adopt the default setting where each sample consists of five snippets of web pages as documents. The experimental results are present in Table 5. In general, we observe a large variance of summary lengths across models. For FSP, ChatGPT and GPT-4 outperform all open-source LLMs on both claim F1 scores and citation F1 scores. However, even GPT-4 exhibits unignorable attribution errors: only 72% of its citations are correct and only 71% of supporting documents are cited. Moreover, only 76% of its generated sentences are fully supported by their respective citations, and only 81% of them are grounded in the input context. For SFT, we observe that smaller pre-trained models such as mT5 significantly lag behind open-source LLMs on both summarization utility and citation quality. Although the fine-tuned mT5-Large model achieves the best groundedness (reflected by the highest ACS score of 90.3%), we find it is primarily due to the model’s tendency to copy the input text rather than summarize the content pertinent to the query, which leads to the suboptimal claim F1 scores. Our additional findings include:
Groundedness errors and citation errors coexist across models.
No model achieves a perfect ACS score, indicating the presence of groundedness errors. Moreover, all models’ ACS scores exceed the respective AIS scores. Since their gaps are simply brought by citation errors, this finding reveals that even if the model grounds its generation contextually, it could still struggle with accurate citations. This underscores the challenge of explicit attribution in both open-source and proprietary LLMs.
Supervised fine-tuning improves attribution.
Without fine-tuning, all open-source LLMs struggle with accurate citations. However, SFT consistently boosts all attribution metrics and narrows the gaps between the AIS and ACS scores, indicating a simultaneous optimization of both groundedness and citation quality. This finding highlights the potential benefits of involving the AQFS task during instruction-tuning (Zhang et al., 2023a) to enhance attribution in open-source LLMs.
Source
Type
Max Doc
Length
# Docs
Claim
F1
Citation
F1
AIS ACS
ChatGPT
Snippet 250 5 54.9 67.8 75.1 84.7
Full Content 512 11 45.9 58.2 72.2 88.9
Full Content 256 20 44.9 51.3 65.2 86.5
ChatGLM3-6B (SFT)
Snippet 250 5 52.3 76.1 81.3 89.4
Full Content 512 11 43.6 51.9 65.8 88.3
Full Content 256 20 41.8 42.9 56.3 84.9
Baichuan2-7B (SFT)
Snippet 250 5 52.8 67.9 71.6 81.7
Full Content 512 11 41.3 51.3 63.7 83.2
Full Content 256 20 41.1 42.8 53.9 80.5
Table 6: Impact of different document settings. Besides the default setting in Table 5, we further evaluate the model performance in summarizing the full content of web pages. We also adopt different chunk sizes (512 and 256) to analyze the impact of document granularity.
5.2 Results in the Long-Context Setting
We further adopt a more challenging long-context setting, where the models are provided with the full content of web pages to summarize. We chunk the web pages into passages with a maximum length of 512 or 256 and assign a unique citation number to each of them. [^10] Our evaluator is based on the mT5 model which supports a context window of 512 tokens. Table 6 presents the performance of the selected models, which shows that:
Figure 4: Performance change over context length of the models in Table 6, where full content of web pages are chunked into documents with a maximum length of 512. Model names are followed by their context window size. The number of input tokens is counted using the tokenizer of each model respectively.
Extending context length reduces model performance.
We observe an overall decline in both summarization utility and attribution with full content instead of snippets. We also visualize the performance variance over the context length in Figure 4, which shows a decline in claim F1, citation F1, and AIS scores as the context length extends. This pattern indicates that extending context length challenges models’ ability to synthesize pertinent information and correctly cite sources.
Using fine-grained documents poses challenges in attribution.
In practice, if the cited documents are too long, it is challenging to verify the model generations, and we expect the models could cite more specific segments of information. Therefore, we also vary the maximum document length when chunking the web pages and investigate the impact of document (citation) granularity. As shown in Table 6, reducing max document length from 512 to 256, without changing the total input evidence, drastically lowers citation F1 and AIS scores for all three models. This performance reduction reveals the difficulty LLMs face in precisely pinpointing the exact supporting evidence within the context.
6 Related Work
Relevant datasets and benchmarks.
Query-focused multi-document summarization (QF-MDS) aims to summarize multiple documents driven by specific queries (Tombros and Sanderson, 1998; Li and Li, 2013; Roy and Kundu, 2023). Similarly, long-form question answering (LFQA) focuses on producing detailed answers, often utilizing external sources (Krishna et al., 2021). Most existing datasets in both tasks do not consider attribution in the setup (Fan et al., 2019; Fabbri et al., 2019; Boni et al., 2021; Stelmakh et al., 2022; Bolotova et al., 2023). Bohnet et al. (2022) propose the attributed question answering (AQA) benchmark where the system must output the answer alongside a piece of evidence. However, long-form responses usually require citing multiple sources of evidence. Recent initiatives (Qin et al., 2023; Liu et al., 2023b; Gao et al., 2023b) in this dimension exhibit limitations in citation annotation and evaluation methods detailed in Section 1 and Table 1.
Evaluating attribution in LLMs.
Attribution refers to the ability to provide external evidence supporting the claims made by the model (Rashkin et al., 2023; Li et al., 2023a). It is crucial for enhancing the credibility of generations and reducing hallucinations (Ji et al., 2023; Zhang et al., 2023c). Although attribution can be approached through various methods, such as generating references from the model’s internal knowledge (Weller et al., 2023), retrieval-augmented generation (RAG) with citations (Nakano et al., 2021; Qin et al., 2023; Liu et al., 2023b; Gao et al., 2023b; Li et al., 2023b), and seeking references post-generation Gao et al. (2023a); Chen et al. (2023); Huo et al. (2023), cost-effective evaluation methods remain a challenging task. Existing automatic metrics Gao et al. (2023b); Bohnet et al. (2022); Yue et al. (2023) solely depend on off-the-shelf NLI models, failing to detect partial support over complex claims. On the other hand, recent works on summarization evaluation Liu et al. (2023c), textual entailment Kamoi et al. (2023), and fact verification Chen et al. (2023); Min et al. (2023) leverage claim decomposition strategy for fine-grained verification. However, they heavily rely on either human efforts or proprietary LLMs to extract sub-claims, which restricts their scalability. Moreover, existing works do not distinguish the evaluation of groundedness and citation quality. This limits the in-depth understanding of attribution errors in different models.
7 Conclusion
In this work, we formulate the task of attributed query-focused summarization (AQFS) and present WebCiteS, a high-quality dataset derived from real-world user queries and search results. We propose a comprehensive evaluation framework on summarization utility and attribution. Notably, we highlight two fine-grained dimensions of attribution: groundedness and citation quality. We further enhance the framework with a carefully-designed automatic evaluator, and validate its substantial agreement with human annotators. Finally, we evaluate both open-source and proprietary LLMs extensively on WebCiteS to underscore the unsolved challenge of attribution, especially in the long-context setting with fine-grained documents. We believe WebCiteS could facilitate more future explorations in attributed language models.
Limitations
Task Setup and Dataset.
The AQFS task and the WebCiteS dataset primarily focus on evaluating and improving the model’s ability to synthesize information from multiple sources with accurate attribution. Therefore, we do not incorporate retrieval into our task setup. Though we find most web pages returned by the search engine are relevant to the queries (Section 2.3), other works highlight the importance of retrieval quality (Gao et al., 2023b; Liu et al., 2023b). We believe the search results and snippets provided in WebCiteS can also serve as a valuable resource to refine open-source retrievers in future works.
Evaluation.
The reliability of our automatic evaluator is dependent on the accuracy of both the claim-split model and the NLI model. Though we have validated their performance in Section 4, we could not ensure the model-generated sub-claims are atomic in granularity. Failing to divide sub-claims could affect the identification of partial support instances. Moreover, the context window of the NLI model is also a constraint. Though the mT5 model uses relative position embeddings Shaw et al. (2018) and accepts arbitrary input sequence length, we find its accuracy drops if the input sequence length significantly exceeds its context window of 512 tokens, primarily due to the length distribution of its training data. Therefore, we believe training a more reliable NLI model for the long-context setting is also an important future work.
Ethics Statement
Since WebCiteS is built upon real user queries, we have taken strict measures to address privacy issues. We only sampled anonymized queries from the search log without collecting any other information such as user identifiers. All queries were shuffled and not present in time order, making it impossible to obtain individual search history from the dataset. Lastly, during annotation, we asked annotators and quality inspectors to pay attention to discard any query with potential privacy issues.
Moreover, we have endeavored to eliminate all inappropriate content from our corpus. Firstly, we adopt an internal commercial tool to automatically detect and discard queries with improper intentions. Secondly, the commercial search engine used in this work has taken content quality and safety into account during web page retrieval and ranking, so it is unlikely that the top five search results would contain dubious or harmful material. Thirdly, human annotators were asked to discard any samples with inappropriate content (see Appendix A). Such a manual inspection process, despite being essential to enhance the dataset quality, would inevitably expose potential risks to the annotators themselves. We attempted to minimize those risks by automatically filtering most inappropriate content with the commercial tool and search engine before annotation. In practice, of all samples discarded by annotators, only less than 5% fell into the category of inappropriate content. Meanwhile, we alerted annotators of the potential risks in advance by including a cautionary note in the annotation instruction and allowed them to skip any sample that made them uncomfortable.
Finally, this work focuses on attribution rather than factuality: even if a response is fully supported by the evidence it attributes, it is not guaranteed to be factual since the evidence itself might contain factual errors or become outdated. Just as previous works (Bohnet et al., 2022; Rashkin et al., 2023) point out, the judge of factuality, especially in open domains, is extremely difficult. During the construction of WebCiteS, though we requested annotators to discard samples with highly questionable materials, we did not assume that their professional expertise could cover all fields in the corpus. Therefore, we emphasize that future works should be cautious about treating the annotated summaries in WebCiteS as "facts".
Appendix A The WebCiteS Dataset
Table 10 presents an annotated example in WebCiteS. Table 7 displays the domain distribution of the user queries. Figure 5 displays the distributions of citation numbers. The distribution of context length for using snippets or full content as documents are shown in Figure 6 and Figure 7 respectively.
Domain Count Domain Count
生活知识 Daily Life Knowledge 1370 金融 Finance 351
教育培训 Education and Training 1032 房产装修 Real Estate and Decoration 256
政策法规 Policies and Regulations 726 生产制造 Manufacturing 221
商品 Commodities 682 游戏娱乐 Gaming and Entertainment 182
其他 Others 508 交通出行 Transportation 158
机动车 Vehicles 430 旅游 Travel 146
信息技术 Information Technology 424 母婴育儿 Maternity and Childcare 135
动植物 Flora and Fauna 398 民俗文化 Folk Culture 111
Table 7: The domain distribution of the user queries in WebCiteS, covering a broad range of real-world scenarios.
Figure 5: The distribution of the number of citations per sentence and summary in WebCiteS.
Figure 6: The distribution of the input context length and the number of input tokens in the default setting, where each sample consists of five snippets of web pages as documents. We use the tokenizer of gpt-3.5-turbo.
Figure 7: The distribution of the input context length and the number of input tokens in the long-context setting, where we chunk the full content of web pages into documents with a maximum length of 512 characters. We use the tokenizer of gpt-3.5-turbo.
A.1 More Details of Data Annotation
Sample selection.
After 40,000 raw queries were gathered, we adopted a rule-based system to remove common trivial queries. We also filter out queries seeking health and medicine advice, since these scenarios are of high risks and hard to judge without professional expertise. Through this process, we obtain 18,500 filtered queries. To ensure the quality of the data, the remaining samples were further filtered by human annotators. Specifically, a sample would be discarded if it matched the following scenarios:
  1. 1.
    If the query was too trivial and did not need long-form answers, or if it was seeking creative inspirations which did not need to be supported by evidence.
  2. 2.
    If the query did not express its demand clearly and was hard to understand.
  3. 3.
    If the query and documents contained inappropriate content, such as personal information, prejudice or bias against specific groups, and controversial topics.
  4. 4.
    If the query could not be answered by the retrieved documents, or the reliability of certain documents was highly questionable to the annotators.
Stage 1: manual screening and information extraction.
We developed annotation software which allowed annotators to highlight clause-level segments containing useful information. The manual filtering of invalid samples based on the above criteria also took place in this stage.
Stage 2: LLM-based candidate summary generation.
In the early annotation phase, We employed ChatGPT to summarize the information extracted from each sample. As our dataset grew, after accumulating 1.2k samples, we additionally fine-tuned an open-source model, ChatGLM2-6B, to provide an extra candidate summary for each sample. We upgraded this model iteratively with the influx of new annotations. Instead of generating multiple outputs by ChatGPT, the motivation for fine-tuning an extra model is to increase the diversity of the candidates.
Stage 3: manual refinement and citation annotation.
We outline a streamlined refinement process as follows: first, annotators were instructed to examine each sentence in the chosen summary, removing unimportant or redundant content. The importance of content was based on the extracted information from the first stage. After that, annotators would identify the verification-worthy sentences in the summary, compare them with all documents with highlighted extraction, and cite all supporting ones. They would also rectify any hallucinations or groundedness errors in the sentences detected during citation annotation. After adding citations, they further ensured all the extracted information was referenced by the summary. If any important information was missing, they would either expand existing sentences or craft new ones to supplement the information, thereby enriching the comprehensiveness of the summary. Finally, annotators inspected the entire summary once again and refined its writing to improve coherence. They were also encouraged to add an introductory sentence to the beginning of the summary to enhance its readability.
Appendix B Evaluation metrics
We bring more details and discussions on the evaluation metrics.
B.1 Metrics of Summarization Utility
Length.
It is measured in the number of characters. We remove all citations in the summary before computing length.
Self-BLEU.
We compute Self-BLEU based on BLEU-4. Our implementation of self-BLEU is based on the sacreBLEU library (Post, 2018).
Claim F1.
Many prior works in summarization evaluation follow the Pyramid protocol (Nenkova and Passonneau, 2004) to decompose the reference summaries into summary content units (Passonneau, 2009; Shapira et al., 2019; Zhang and Bansal, 2021) or atomic content units (Liu et al., 2023c), and measure how many units are covered in the system summaries (i.e., recall-based metrics). One advantage of recall-based metrics is that only reference summaries need to be decomposed since early works mostly rely on human efforts for sentence decomposition. On the other hand, recent studies also investigate the use of proprietary LLMs Gao et al. (2023b); Min et al. (2023); Kamoi et al. (2023). For example, Gao et al. (2023b) use InstructGPT Ouyang et al. (2022) to generate three sub-claims for each reference answer and compute claim recall to measure the correctness of model generations. However, the cost and rate limits of proprietary LLMs still hinder the scalability. In this work, we train a tailored claim-split model that enables us to calculate both claim precision and recall by extracting all sub-claims from both the system summaries and reference summaries with minimum cost. Moreover, we do not use traditional automatic metrics such as ROUGE (Lin, 2004) since their limitations in evaluating LLM generations have been discussed in recent works. (Zhang et al., 2023b; Gao et al., 2023a; Xu et al., 2023).
B.2 Metrics of Attribution
Granularity.
We evaluate all metrics of attribution at the sentence level, primarily because the citations in the WebCiteS are annotated at the sentence level, as more fine-grained annotation would require annotators to manually extract sub-claims from a sentence which bring extra cost. However, future works could extend these metrics to the sub-claim level depending on their needs.
Citation precision.
Liu et al. (2023a); Gao et al. (2023b) compute citation precision by calculating the fraction of accurate citations within the whole generation. The major differences in our approach are:
  1. 1.
    We try to look for $C^{i*}_{\text{Pred}}$ if $C^{i}_{\text{Pred}}$ is empty to avoid unnecessary penalization which leads to the undervaluation of model performance.
  2. 2.
    We compute citation precision at the sentence level and average them for the whole summary, while Liu et al. (2023a); Gao et al. (2023b) compute this metric directly at the response-level.
Citation recall.
Liu et al. (2023a); Gao et al. (2023b) define citation recall as the fraction of sentences being fully supported by their citations. This is equivalent to the AIS score (Bohnet et al., 2022; Rashkin et al., 2023; Gao et al., 2023a) by taking citations as the identified sources. In contrast, our definition of citation recall is consistent with citation precision by calculating the fraction of citations, which is aligned with the naming of the metric.
Citation Mask
Default Auto Human
XLM-RoBERTa-Large-XNLI 75.1 78.0 78.4
ELS-RoBERTa-Large-NLI 74.2 76.2 77.6
ELS-MBERT-1.3B-NLI 76.0 79.1 79.3
mT5-Large-XNLI 78.6 82.3 82.6
Table 8: Performance of citation prediction using different NLI models under three citation mask settings.
Appendix C Experiments on NLI Model
We evaluate the performance of different NLI models via a citation prediction task on the test set of WebCiteS: for each sentence in the summary and each given document, we use the NLI model to classify whether the sentence should cite the document, and calculate its accuracy by taking human citations as ground truth. Only sentences with citation mask $m_{i}=1$ are considered. We adopt three citation mask settings: default, auto, and human, similar to the experiments in Section 4.3. We select the following NLI models for evaluation: (1) XLM-RoBERTa-Large-XNLI, an XLM-RoBERTa model (Conneau et al., 2020) fine-tuned on multilingual NLI datasets. [^11] https://huggingface.co/joeddav/xlm-roberta-large-xnli (2) ELS-RoBERTa-Large-NLI, a Chinese RoBERTa model fine-tuned on several NLI datasets (Zhang et al., 2022). [^12] https://huggingface.co/IDEA-CCNL/Erlangshen-Roberta-330M-NLI (3) ELS-MBERT-1.3B-NLI, a Chinese model based on the MegatronBERT architecture (Shoeybi et al., 2019), fine-tuned on several NLI datasets (Zhang et al., 2022). [^13] https://huggingface.co/IDEA-CCNL/Erlangshen-MegatronBert-1.3B-NLI, (4) mT5-Large-XNLI, an mT5 model (Xue et al., 2021) fine-tuned on multilingual NLI datasets. [^14] https://huggingface.co/alan-turing-institute/mt5-large-finetuned-mnli-xtreme-xnli
The results in Table 8 show that the mT5 model achieves the highest accuracy in predicting citations. Moreover, using the default citation mask (i.e., set $m_{i}=1$ for all sentences) lowers accuracy across all models, underscoring the necessity of identifying if a sentence is verification-worthy. Besides, we find that results under auto citation mask and human citation mask are notably similar. This validates the effectiveness of our citation mask prediction method proposed in Section 3.2.
Claim Citation
Source Type Max Doc Length # Docs Len. Self- Bleu ↓ Prec. Rec. F1 Prec. Rec. F1 AIS ACS
ChatGPT
Snippet 250 5 223 12.2 53.5 56.5 54.9 71.2 64.8 67.8 75.1 84.7
Full Content 512 11 238 10.5 42.9 49.3 45.9 63.6 53.6 58.2 72.2 88.9
Full Content 256 20 245 10.6 41.2 49.3 44.9 58.2 45.9 51.3 65.2 86.5
ChatGLM3-6B (SFT)
Snippet 250 5 163 8.9 55.7 49.3 52.3 78.5 73.8 76.1 81.3 89.4
Full Content 512 11 170 8.9 45.6 40.8 43.1 61.7 44.7 51.9 65.8 88.3
Full Content 256 20 173 8.9 44.5 39.4 41.8 53.1 35.9 42.9 56.3 84.9
Baichuan2-7B (SFT)
Snippet 250 5 212 9.9 52.4 53.1 52.8 68.1 67.8 67.9 71.6 81.7
Full Content 512 11 208 10.4 42.0 40.7 41.3 58.9 45.4 51.3 63.7 83.2
Full Content 256 20 208 10.4 41.3 40.9 41.1 50.4 37.3 42.8 53.9 80.5
Table 9: Full results of model performance in different document settings shown in Table 6.
Appendix D Experiments on Claim-Split Model
Data for training and evaluation.
As described in Section 4.2, our approach involves fine-tuning mT5 models with ChatGPT outputs. We craft a comprehensive prompt with detailed instructions, as shown in Table 13. With ChatGPT’s feature of structuring JSON outputs, we prompt it to extract all the sub-claims from sentences within the entire summary in a single response, and then split the response into sentence-level claim-split outputs for training and evaluation. We divide the outputs into training, validation, and test sets aligning with the split of the sample in the WebCiteS dataset. However, since the granularity of sub-claims generated by ChatGPT is not always atomic, we make additional adjustments to the data distribution: we keep all sentences with more than one sub-claim and sample an equal number of sentences with a single sub-claim (either because they are not divisible or because ChatGPT fails to separate their sub-claims). This results in a distribution of 16,158 sentences for training, 2,858 for development, and 1,330 for testing.
Implementation details.
For fine-tuning, we use the batch size of 64 and the learning rate of 1e-4. We use the AdamW (Loshchilov and Hutter, 2019) optimizer and train the models for 5 epochs. For inference, we use greedy decoding and load the model in half-precision to accelerate evaluation.
Appendix E Experiments on the AQFS Task.
E.1 Implementation Details
Few-shot prompting (FSP).
Utilizing the in-context learning abilities of LLMs Brown et al. (2020), we construct a prompt with four parts:
  • Instruction: A paragraph that introduces the task and describes specific requirements.
  • Demonstration: An example with the query, source documents, and human-annotated summary as reference.
  • Sample to Summarize: The query and source documents that the model needs to summarize.
  • Ending: An ending statement guiding the model to produce the summary as required.
The full prompt is displayed in Table 10.
Supervised fine-tuning (SFT).
we also fine-tune open-source models in our experiments. To save GPU memory, we condense the above prompt as the input text by shortening the instruction and removing the demonstration. The condensed instruction is present in Table 11.
For mT5 models, we use the batch size of 64, the learning rate of 1e-4, AdamW as the optimizer, and fine-tune them for 5 epochs. For other open-source LLMs, we use the same batch size and optimizer, while setting the learning rate to 2e-5 and fine-tuning them for 1 epoch, as we find more epochs lead to the rise of validation loss. All open-source LLMs are trained on 8 NVIDIA A100 40G GPUs using Deepspeed ZeRO Stage-3 framework(Rajbhandari et al., 2020). We adopt FP16 mixed precision training (Micikevicius et al., 2018) for ChatGLM2 and ChatGLM3, and BF16 mixed precision training for Baichuan2 models, based on their default configurations.
Inference.
For ChatGPT and GPT-4, we use the default parameters of the OpenAI Chat API; for open-source models, we follow Gao et al. (2023b) to use Nucleus sampling Holtzman et al. (2020) with top_p=0.95. We load open-source LLMs in either FP16 or BF16 precision to accelerate inference and save GPU memory.
Data for the long context setting.
In Section 5.2, we adopt a long-context setting where the models are provided with the full content of web pages to summarize. We chunk the web page into documents with a maximum length of 512 or 256 characters. The chunking is performed at the sentence level, where we try to avoid splitting a single sentence into multiple documents. Web pages shorter than the maximum document length are directly taken as the documents.
Table 10: The prompt we use for few-shot prompting with full instruction. Table 12 presents the translation.
Table 11: The condensed instruction used for supervised fine-tuning. The translation is in italic text.
Table 12: The translation of the prompt in Table 10.
Table 13: The prompt we used to split sub-claims with ChatGPT. The translation is in italic text.